# Iñigo Alonso1, Imanol Miranda2, Eneko Agirre2, Mirella Lapata1

### v2

@@ Abstract
While table understanding increasingly relies on pixel-only settings where tables are processed as visual representations, current benchmarks predominantly use synthetic renderings that lack the complexity and visual diversity of real-world tables. Additionally, existing visual table understanding (VTU) datasets offer fixed examples with single visualizations and pre-defined instructions, providing no access to underlying serialized data for reformulation. We introduce Tablet, a large-scale VTU dataset with 4 million examples across 20 tasks, grounded in 2 million unique tables where 88% preserve original visualizations. Each example includes paired image-HTML representations, comprehensive metadata, and provenance information linking back to the source datasets. Fine-tuning vision-language models like Qwen2.5-VL-7B on Tablet improves performance on seen and unseen VTU tasks while increasing robustness on real-world table visualizations. By preserving original visualizations and maintaining example traceability in a unified large-scale collection, Tablet establishes a foundation for robust training and extensible evaluation of future VTU models. [^1] Dataset, code, and other resources are available at https://github.com/alonsoapp/TABLET
@@ 1 Introduction
The field of table understanding focuses on techniques for representing and interpreting tabular data to support a wide range of practical tasks such as question answering, summarization, and information extraction. Research in this area has traditionally representated tables as structured text, encoding their content and layout through linearized or graph-based representations (see Figure 1b; herzig-etal-2020-tapas; zhang2020graph; liu2022tapex). While this unimodal view remains effective in certain domains, many tables found in documents and webpages contain irregular structures, rely on visual formatting (e.g., merged cells, background colors, font variations), or embed multimodal elements such as images (see Figure 1a). Advances in Vision-Language Models (VLMs; pmlr-v139-radford21a; liu2023llava) have provided impetus for treating tables as images, eschewing the step of rendering them as text sequences (like Markdown or HTML). The conceptual simplicity of this approach, coupled with improved performance on several tabular tasks (alonso-etal-2024-pixt3; zhou-etal-2025-texts) has driven significant research interest (zheng-etal-2024-multimodal; su2024tablegpt2largemultimodalmodel; jiang2025multimodaltabularreasoningprivileged) in Visual Table Understanding (also known as Multimodal Table Understanding). Visual representations of tables are not only merely convenient but in many cases necessary, particularly for VLM agents that interact with the world exclusively through pixels (e.g., on a screen) and must interpret tables directly in their visual form (deng2023mind2web; zheng2024gptvision; lu2024omniparserpurevisionbased).
@@
Despite the growing relevance of VTU, there are few resources that support training models directly on image-based representations of tables. Existing benchmarks like MMTab (zheng-etal-2024-multimodal) consist of web tables (e.g., from Wikipedia which is a common source for many tabular datasets), that are serialized and subsequently rendered as synthetic images (see Figures 1b,c). These images do not preserve the original visual characteristics of the tables and do not reflect the diversity and complexity of real-world layouts. As a result, models trained on such data face a train-test mismatch, since the visual patterns learned from serialized renderings do not generalize well to naturally occurring tables failing to capture critical visual cues like subtle ruling lines, intricate merged cell layouts, background colors, font variations, or embedded images that are inherent to real-world table comprehension (compare Figure 1a and 1c). An exception is WikiDT shi2024wikidT, which provides access to original table visualizations from web pages but focuses exclusively on table question answering. Likewise, datasets designed for processing PDF documents or screenshots (zhong2020image; lu2023dynamic) preserve original table visualizations but support only a single task.
In this work, we introduce Tablet [^2] Our dataset is called Tablet partly as a nod to the Scottish sugary confection which is often cut into uniform squares or rectangles, resembling a table., a large-scale dataset designed to enable vision-language models to learn generalizable skills for table understanding by leveraging visual features from table images. Unlike existing datasets that focus on a single task or use synthetic renderings, Tablet preserves lossless representations of real-world table images and spans a diverse set of tasks, including table-to-text generation (parikh-etal-2020-totto), table fact verification (chen2020hybridqa), and table-based question answering (pasupat-liang-2015-compositional). Tablet facilitates training and evaluation on table formats that align with those encountered in downstream applications, particularly in pixel-based agents that process visual input directly. It contains 4 million examples across 20 tasks, derived from 2 million unique tables, 88% of which retain their original visualization. To promote flexibility and foster future research, we provide both image and HTML representations of tables, together with metadata and links to the source datasets. Pairing table images with HTML enables models to learn from naturally occurring and synthetic table images when the former are not available. Tablet’s large-scale and task aggregation provide a unique, foundational resource that complements research integrating textual and visual modalities for training and reasoning (liu2025hippoenhancingtableunderstanding; jiang2025multimodaltabularreasoningprivileged). We experimentally demonstrate that fine-tuning on Tablet improves performance on in- and out-of-domain benchmarks. Our work makes the following contributions:
  • Tablet addresses a critical gap in large-scale, visually faithful datasets for VTU, with 4 million examples and 2 million tables, most of which are preserved in their original form.
  • We show that supervised fine-tuning on Tablet  enhances VLM performance across seen and unseen table understanding tasks. We release a VLM model trained on Tablet that achieves state-of-the-art results on 10 out of 14 table understanding benchmarks.
  • Tablet  is designed for extensibility: all tasks are provided in a unified format, paired with HTML table serializations that can be rendered back into the original visualization, along with example metadata and links to the source datasets to facilitate reuse and reformulation.
@@ 2 Related Work
Early work approached the problem of table understanding from a unimodal perspective, treating tables as structured text and developing general-purpose models that generalized across multiple tasks (li2023tablegpttabletunedgptdiverse; zhang-etal-2024-tablellama). Recent progress in natural image understanding afforded by increasingly better VLMs (pmlr-v139-radford21a; liu2023llava) has led to development of models that can successfully process visually rendered text beyond natural imagery and perform table understanding holistically, within a multimodal framework (kim2022donut; lee2023pix2structscreenshotparsingpretraining; zhang2023llavar; ye-etal-2023-ureader; hu-etal-2024-mplug; alonso-etal-2024-pixt3; jiang2025multimodaltabularreasoningprivileged; su2024tablegpt2largemultimodalmodel; zhao2024tabpedia; zheng-etal-2024-multimodal; su2024tablegpt2largemultimodalmodel).
To this effect, several benchmarks have been proposed recently to evaluate progress in VTU. For instance, TableVQA-Bench (kim2024tablevqabenchvisualquestionanswering) focuses on visual table question answering but relies on synthetic table images. In contrast, MMTBench (titiya2025mmtbenchunifiedbenchmarkcomplex) includes original table visualizations, along with visually rich images and interleaved charts. While both are valuable for evaluation, they are limited in scope and do not provide large-scale training data, restricting their utility for developing generalizable models.
Larger training datasets (with training splits) have historically focused on image-based Table Structure Recognition (TSR), including PubTabNet (zhong2020image), TableBank (li-etal-2020-tablebank), and TabComp (gautam-etal-2025-tabcomp). Beyond TSR, alonso-etal-2024-pixt3 created image-rendered versions of existing datasets like ToTTo (parikh-etal-2020-totto), relying on lossy renderings that discard visual features. WikiDT (shi2024wikidT) preserves original table visualizations but only targets Visual TableQA. MMTab (zheng-etal-2024-multimodal) is a large VTU dataset, with 150k TSR pre-training examples and 232k instruction examples across 19 tasks. While MMTab is a useful resource that we build upon, it relies on synthetic renderings of serialized tables, lacks traceability to original data sources, and remains limited in size compared to the scale needed for general-purpose VTU.
In this work, we do not introduce a task-specific dataset or model. Instead, we create a large-scale resource aimed at enhancing table understanding in general-purpose vision-language models like Gemma-3 (gemmateam2025gemma3technicalreport) or Qwen2-VL (qwen2vl), under the assumption that tables for many tasks are seen as images, and therefore can be naturally processed by these models. This distinguishes Tablet from many recent efforts that introduce task-specific datasets or specialized architectures as it is designed to be a holist resource for advancing VLM capabilities for tables.
@@ 3.1 Overview
Tablet is a large-scale dataset that aggregates a total of 4,066,545 examples, combining 20 TU tasks sourced from 14 datasets (see Section 3.3 for a breakdown of examples per task). Tablet includes 2,031,256 unique table images with visualizations from Wikipedia (61.5%), PubTabNet (25.1%; zhong2020image), TabMWP (1.9%; lu2023dynamic, and synthetically rendered images (11.5%) from serialized tables (see Section 3.2 for details).
Examples in Tablet are framed as instructions, leveraging the benefits of instruction-tuning for large language models.Tablet does not introduce new data instances; rather it repurposes examples from existing datasets, referred to as seed datasets, rephrasing their tasks into an instruction format. Examples are drawn from 14 seed datasets: TURL (deng2020turl), ToTTo (parikh-etal-2020-totto), TabFact (2019TabFactA), WikiTableQuestions (pasupat-liang-2015-compositional), HybridQA (chen2020hybridqa), HiTab (cheng-etal-2022-hitab), PubTabNet (zhong2020image), TabMWP (lu2023dynamic), TAT-QA (zhu-etal-2021-tat), InfoTabs (gupta-etal-2020-infotabs), WikiBIO (lebret-etal-2016-neural), FeTaQA (nan-etal-2022-fetaqa), MMTab zheng-etal-2024-multimodal, and DocStruct4M (hu-etal-2024-mplug).
Datasets that include only images and fixed prompts restrict researchers to the provided visualizations and instruction formulations, limiting the exploration of new prompting techniques or the use of table data to create additional examples. To avoid this, each example in Tablet is provided in a unified format, including the instruction used in this work, its corresponding atomic data, the HTML version of the table, the original source ID, and additional metadata (detailed in Appendix7.4).
@@
Total 3,505,311 210,041 351,193
Table 1: Tablet splits, tasks, and seed datasets; M1: InfoTabs, NSF, StatCan, TabMWP, TAT-QA, TabFact, ToTTo, WikiBIO, WikiTableQuestions; M2: InfoTabs, NSF, StatCan, TAT-QA, TabFact, ToTTo, WikiBIO; M3: PubTabNet, TabFact, WikiTableQuestions. Proportion of examples in each set (train/dev/test) is shown in gray within parentheses.
@@ 3.2 Image sources
Wikipedia Tables Wikipedia tables generally follow a hierarchical table structure and are often rich in visual information. 61.5% of the tables in Tablet are lossless visualizations sourced from Wikipedia (See Appendix 7.2 for a breakdown of image sources per task). These tables are referenced in 67.4% of all examples in the dataset. Each table from the seed datasets was traced back to its original visualization in the corresponding Wikipedia article and captured in a screenshot. [^3] Table images were rendered using Firefox in headless mode (version 142.0.1 with GeckoDriver version 0.36.0) at an effective rendering density of 96 PPI (which stands for Pixels Per Inch).. A key challenge in retrieving the original visualizations was that the seed datasets were created at different times, while Wikipedia articles are continuously updated. To address this, we relied on the metadata released with the seed datasets and also reached out to their authors to find out when the tables were harvested (see Appendix 7.3). We used Wikipedia’s archiving API to retrieve each article as it appeared at the time of crawling. All Wikipedia tables in Tablet are linked to their source via both the page identifier and the corresponding revision identifier (oldid). As Wikipedia pages often contain multiple tables, we needed to identify which one corresponded to the seed example. We did this by computing the levenshtein1966binary edit distance between Wikipedia tables and seed tables, both represented in markdown format, and selecting those with the highest similarity value. We set a minimum similarity threshold of 0.70; in cases without a match, the serialized seed table was rendered as an image.
Synthetic Tables Wikipedia tables that could not be retrieved, either due to low similarity scores or other issues, were synthetically generated by converting their serialized format to HTML and rendering them using the same browser configuration used for the original Wikipedia tables. These synthetic tables account for 11.5% of the dataset, contributing 747,002 additional examples (18.4% of the total). In Section 5, we show that combining original table images with synthetic visualizations during training allows for an increase in example count and visualization diversity that leads to better model performance compared to using synthetic or original tables alone.
Document Tables We also incorporate tables from other domains to improve model generalization. These include tables from scientific articles in PubTabNet (zhong2020image) and rendered tables from TabMWP (lu2023dynamic). PubTabNet contributes 509,892 tables to our dataset (25.1% of the total), accounting for 97.5% of the structure-aware learning task incorporated from DocStruct4M (hu-etal-2024-mplug). TabMWP adds 38,431 tables (1.9% of the total); while smaller in scale, TabMWP offers valuable coverage of numerical reasoning examples.
@@ 3.3 Table Understanding Tasks
Tablet aggregates a total of 20 TU tasks grouped into 7 broad categories: Table Interpretation, Question Answering, Table-to-Text Generation, Table Numerical Reasoning, Table Natural Language Inference (NLI), and Table Structure Understanding. Subtasks within each category require different skills and demonstrate various degrees of complexity, ranging from basic table structure understanding to downstream tasks that combine tabular reasoning with other skills, such as information retrieval, numerical reasoning, or natural language inference. Table 1 provides a breakdown of examples per task and their source dataset. We briefly describe each task below and provide more details in Appendix 7.5.
@@
Table Interpretation comprises three tasks, namely Column Type Annotation, Entity Linking, and Relation Extraction (ent_link, col_type, and rel_extraction in Table 1). Sourced from TURL (deng2020turl), these tasks do not represent downstream applications but target basic table understanding skills essential for tackling more complex tasks. All instructions for these tasks were aggregated from TURL. They constitute the largest group in our training set, with 1,899,149 examples (54.2%).
@@
Table Question Answering is represented by Free-form Table QA (FeTaQA, nan-etal-2022-fetaqa), Hierarchical Table QA (HiTabQA; cheng-etal-2022-hitab), Hybrid QA (HybridQA; chen2020hybridqa), and Table QA (WikiTableQuestions; pasupat-liang-2015-compositional). These tasks range from simple question answering over table content to multi-hop reasoning that combines textual and tabular information. They also vary in terms of table complexity and the expected answer format. Question answering represents 2.5% of Tablet’s training set (87,245 examples).
@@
Table-to-Text Generation is exemplified by two subtasks, namely Highlighted Cell Text Generation (ToTTo; parikh-etal-2020-totto) and Wikipedia Biography Generation (WikiBio; lebret-etal-2016-neural). Both tasks involve generating text based on table content, concise biographies in the case of WikiBio, and short descriptions based on visually highlighted cells, in the case of ToTTo. Together, they account for 693,593 training examples (19.8%).
@@
Table Numerical Reasoning combines Tabular Math Word Problem Solving (TabMWP; lu2023dynamic) and Hybrid-context Financial QA (TAT-QA; zhu-etal-2021-tat). Given a table and a mathematical question, the model must perform reasoning over table values. This type of numerical reasoning represents 0.8% of Tablet’s training set (25,260 examples).
@@
Table NLI includes two entailment tasks: Table Fact Verification (TabFact; 2019TabFactA) and Infobox Natural Language Inference (InfoTabs; gupta-etal-2020-infotabs) where statements as supported or refuted based on table content. They represent 104,255 examples (3%) in Tablet’s training set.
@@
Table Structure Understanding is an umbrella category for tasks designed to facilitate structural understanding of tables: Merged Cell Detection, Row and Column Extraction, Table Cell Extraction, Table Cell Location, Table Recognition, Table Size Detection, and Structure-aware Parsing in Table 1). Examples for the first seven tasks are aggregated from MMTab, while Structure-aware Parsing is derived from DocStruct4M (hu-etal-2024-mplug). These tasks are based on tables from multiple seed datasets (see rows with seeds M2, M1, and M3), including InfoTabs (gupta-etal-2020-infotabs), NSF (nsf_science_engineering_indicators_2019), StatCan (statistics_canada_2024), TabMWP (lu2023dynamic), TAT-QA (zhu-etal-2021-tat), TabFact (2019TabFactA), ToTTo (parikh-etal-2020-totto), WikiBio (lebret-etal-2016-neural), WikiTableQuestions (pasupat-liang-2015-compositional), and PubTabNet zhong2020image. While we maintain the original MMTab instructions, we use our own table visualizations. This group contributes 558,865 examples (15.8%) to our training.

### v3

@@ Abstract
While table understanding increasingly relies on pixel-only settings, current benchmarks predominantly use synthetic renderings that lack the complexity and visual diversity of real-world tables. Additionally, existing visual table understanding (VTU) datasets offer fixed examples with single visualizations and pre-defined instructions, providing no access to underlying serialized data for reformulation. We introduce Tablet, a large-scale VTU dataset with 4 million examples across 21 tasks, grounded in 2 million unique tables where 88% preserve original visualizations. To evaluate whether models are able to jointly reason over tabular and visual content, we also introduce VisualTableQA, a benchmark requiring both visual perception and table understanding. Fine-tuning vision-language models like Qwen2.5-VL-7B and Gemma 3-4B on Tablet improves performance on seen and unseen VTU tasks while increasing robustness on real-world table visualizations. By preserving original visualizations and maintaining example traceability in a unified large-scale collection, Tablet establishes a foundation for robust training and extensible evaluation of future VTU models. [^1] Dataset, code, and other resources are available at https://github.com/alonsoapp/TABLET.
@@ 1 Introduction
The field of table understanding focuses on techniques for representing and interpreting tabular data to support a wide range of practical tasks such as question answering, summarization, and information extraction. Research in this area has traditionally represented tables as structured text, encoding their content and layout through linearized or graph-based representations (see Figure 1b; Herzig et al. 2020; Zhang et al. 2020; Liu et al. 2022). While this unimodal view remains effective in certain domains, many tables found in documents and webpages contain irregular structures, rely on visual formatting (e.g., merged cells, background colors, font variations), or embed multimodal elements such as images (see Figure 1a). Advances in Vision-Language Models (VLMs; Radford et al. 2021; Liu et al. 2023) have provided impetus for treating tables as images, eschewing the step of rendering them as text sequences (like Markdown or HTML). The conceptual simplicity of this approach, coupled with improved performance on several tabular tasks (Alonso et al., 2024; Zhou et al., 2025) has driven significant research interest (Zheng et al., 2024b; Su et al., 2024; Jiang et al., 2025) in Visual Table Understanding (also known as Multimodal Table Understanding). Visual representations of tables are not only merely convenient but in many cases necessary, particularly for VLM agents that interact with the world exclusively through pixels (e.g., on a screen) and must interpret tables directly in their visual form (Deng et al., 2023; Zheng et al., 2024a; Lu et al., 2024).
@@
Despite the growing relevance of VTU, there are few resources that support training models directly on image-based representations of tables. Existing benchmarks like MMTab (Zheng et al., 2024b) consist of web tables (e.g., from Wikipedia which is a common source for many tabular datasets), that are serialized and subsequently rendered as synthetic images (see Figures 1b,c). These images do not preserve the original visual characteristics of the tables and do not reflect the diversity and complexity of real-world layouts. As a result, models trained on such data face a train-test mismatch, since the visual patterns learned from serialized renderings do not generalize well to naturally occurring tables failing to capture critical visual cues like subtle ruling lines, intricate merged cell layouts, background colors, font variations, or embedded images that are inherent to real-world table comprehension (compare Figure 1a and 1c). An exception is WikiDT Shi et al. (2024), which provides access to original table visualizations from web pages but focuses exclusively on table question answering. Likewise, datasets designed for processing PDF documents or screenshots (Zhong et al., 2020; Lu et al., 2023) preserve original table visualizations but support a single task.
@@
In this work, we introduce Tablet [^2] Our dataset is called Tablet partly as a nod to the Scottish sugary confection which is often cut into uniform squares or rectangles, resembling a table., a large-scale dataset designed to enable vision-language models to learn generalizable skills for table understanding by leveraging visual features from table images. Unlike existing datasets that focus on a single task or use synthetic renderings, Tablet preserves lossless representations of real-world table images and spans a diverse set of tasks, including table-to-text generation (Parikh et al., 2020), table fact verification (Chen et al., 2020b), and table-based question answering (Pasupat and Liang, 2015). Tablet facilitates training on table formats that align with those encountered in downstream applications, particularly in pixel-based agents that process visual input directly. It contains 4 million examples across 21 tasks, derived from 2 million unique tables, 88% of which retain their original visualization. To promote flexibility and foster future research, we provide both image and HTML representations of tables, together with metadata and links to the source datasets. Pairing table images with HTML enables models to learn from naturally occurring and synthetic table images when the former are not available. Tablet’s large scale and task diversity make it a foundational resource for research on integrating textual and visual modalities (Liu et al., 2025; Jiang et al., 2025).
To demonstrate Tablet’s effectiveness as a training resource, we introduce VisualTableQA, a Visual Table Question Answering benchmark that pairs visually rich tables with questions that cannot be answered through table structure parsing or visual cues alone, but demand joint reasoning over both modalities. Our experiments demonstrate that fine-tuning on Tablet yields substantial improvements on VisualTableQA, as well as strong transfer to other in- and out-of-domain benchmarks. Our work makes the following contributions:
  • We introduce Tablet to address a critical gap in VTU datasets, providing 4 million examples from 2 million tables, most preserved in their original visualization. We design Tablet for extensibility with a unified task format, HTML serializations that render back to original visualizations, and metadata with source links to facilitate reuse.
  • Supervised fine-tuning on Tablet substantially improves VLM performance on both seen and unseen tasks. Our best-performing model achieves state-of-the-art results on 11 out of 15 table understanding benchmarks.
  • We introduce VisualTableQA, a benchmark requiring joint table understanding and visual perception and show that fine-tuning on Tablet improves performance on this challenging unseen task.
@@ 2 Related Work
Early work approached the problem of table understanding from a unimodal perspective, treating tables as structured text and developing general-purpose models that generalized across multiple tasks (Li et al., 2023; Zhang et al., 2024). Recent progress in natural image understanding afforded by increasingly better VLMs (Radford et al., 2021; Liu et al., 2023) has led to development of models that can successfully process visually rendered text beyond natural imagery and perform table understanding holistically, within a multimodal framework (Kim et al., 2022; Lee et al., 2023; Zhang et al., 2023; Ye et al., 2023; Hu et al., 2024; Alonso et al., 2024; Jiang et al., 2025; Su et al., 2024; Zhao et al., 2024; Zheng et al., 2024b; Su et al., 2024).
To this effect, several benchmarks have been proposed recently to evaluate progress in VTU. For instance, TableVQA-Bench (Kim et al., 2024) focuses on visual table question answering but relies on synthetic table images and does not require visual reasoning: its questions can be answered from textual content alone. In contrast, MMTBench (Titiya et al., 2025) includes original table visualizations, along with visually rich images and interleaved charts. While both are valuable for evaluation, they are limited in scope and do not provide large-scale training data, restricting their utility for developing generalizable models.
Larger training datasets have historically focused on image-based Table Structure Recognition (TSR), including PubTabNet (Zhong et al., 2020), TableBank (Li et al., 2020), and TabComp (Gautam et al., 2025). Beyond TSR, Alonso et al. (2024) created image-rendered versions of existing datasets like ToTTo (Parikh et al., 2020), relying on lossy renderings that discard visual features. WikiDT (Shi et al., 2024) preserves original table visualizations but only targets question answering. MMTab (Zheng et al., 2024b) is a large VTU dataset, with 150k TSR pre-training examples and 232k instruction examples across 19 tasks. While MMTab is a useful resource that we build upon, it relies on synthetic renderings of serialized tables, lacks traceability to original data sources, and remains limited in size compared to the scale needed for general-purpose VTU.
In this work, we do not introduce a task-specific dataset or model. Instead, we create a large-scale resource aimed at enhancing table understanding in general-purpose vision-language models like Gemma-3 (Team et al., 2025) or Qwen2-VL (Wang et al., 2024), under the assumption that tables for many tasks are seen as images, and therefore can be naturally processed by these models. This distinguishes Tablet from many recent efforts that introduce task-specific datasets or specialized architectures as it is designed to be a holist resource for advancing VLM capabilities for tables.
@@ 3.1 Overview
Tablet is a large-scale dataset that aggregates a total of 4,066,851 examples, combining 21 TU tasks sourced from 14 datasets (see Section 3.3 for a breakdown of examples per task). Tablet includes 2,031,256 unique table images with visualizations from Wikipedia (61.5%), PubTabNet (25.1%; Zhong et al. 2020), TabMWP (1.9%; Lu et al. 2023, and synthetically rendered images (11.5%) from serialized tables (see Section 3.2 for details).
Examples in Tablet are framed as instructions, leveraging the benefits of instruction-tuning for large language models.Tablet does not introduce new training data instances; rather it repurposes examples from existing datasets, referred to as seed datasets, rephrasing their tasks into an instruction format. Examples are drawn from 14 seed datasets: TURL (Deng et al., 2020), ToTTo (Parikh et al., 2020), TabFact (Chen et al., 2020a), WikiTableQuestions (Pasupat and Liang, 2015), HybridQA (Chen et al., 2020b), HiTab (Cheng et al., 2022), PubTabNet (Zhong et al., 2020), TabMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), InfoTabs (Gupta et al., 2020), WikiBIO (Lebret et al., 2016), FeTaQA (Nan et al., 2022), MMTab Zheng et al. (2024b), and DocStruct4M (Hu et al., 2024). In addition, Tablet includes 306 newly created, human-annotated examples of questions based on visually rich tables drawn from six of the seed datasets.
Datasets that contain only images and fixed prompts restrict researchers to the provided visualizations and instruction formulations, limiting the exploration of new prompting techniques or the use of table data to create additional examples. To avoid this, each example in Tablet is provided in a unified format, including the instruction used in this work, its corresponding atomic data, the HTML version of the table, the original source ID, and additional metadata (detailed in Appendix7.4).
@@
visual_table_qa M4 0 (0.0%) 0 (0.0%) 306 (0.1%)
Total 3,505,311 210,041 351,499
Table 1: Tablet splits, tasks, and seed datasets; M1: InfoTabs, NSF, StatCan, TabMWP, TAT-QA, TabFact, ToTTo, WikiBIO, WikiTableQuestions; M2: InfoTabs, NSF, StatCan, TAT-QA, TabFact, ToTTo, WikiBIO; M3: PubTabNet, TabFact, WikiTableQuestions; M4: ToTTo, HybridQA, TabFact, WikiBIO, TURL, WikiTableQuestions.
@@ 3.2 Image sources
Wikipedia Tables Wikipedia tables generally follow a hierarchical structure and are often rich in visual information. 61.5% of the tables in Tablet are lossless visualizations sourced from Wikipedia (see Appendix 7.2 for a breakdown of image sources per task). These tables are referenced in 67.4% of all examples in the dataset. Each table from the seed datasets was traced back to its original visualization in the corresponding Wikipedia article and captured in a screenshot. [^3] Table images were rendered using Firefox in headless mode (version 142.0.1 with GeckoDriver version 0.36.0) at an effective rendering density of 96 PPI (which stands for Pixels Per Inch).. A key challenge in retrieving the original visualizations was that the seed datasets were created at different times, while Wikipedia articles are continuously updated. To address this, we relied on the metadata released with the seed datasets and also reached out to their authors to find out when the tables were harvested (see Appendix 7.3). We used Wikipedia’s archiving API to retrieve each article as it appeared at the time of crawling. All Wikipedia tables in Tablet are linked to their source via the page identifier and the corresponding revision identifier (oldid). As Wikipedia pages often contain multiple tables, we identified which one corresponded to the seed example by computing the Levenshtein (1966) edit distance between the tables represented in markdown format, and selecting those with the highest similarity value. We set a minimum similarity threshold of 0.7. In cases without a match, the serialized seed table was rendered as an image.
Synthetic Tables Wikipedia tables that could not be retrieved, either due to low similarity scores or other issues, were synthetically generated by converting their serialized format to HTML and rendering them using the same browser configuration used for the original Wikipedia tables. These synthetic tables account for 11.5% of the dataset, contributing 747,002 additional examples (18.4% of the total). In Section 5, we show that combining original table images with synthetic visualizations during training leads to improved performance over using either alone, due to increased example count and visualization diversity.
Document Tables We also incorporate tables from other domains to improve model generalization. These include tables from scientific articles in PubTabNet (Zhong et al., 2020) and rendered tables from TabMWP (Lu et al., 2023). PubTabNet contributes 509,892 tables to our dataset (25.1% of the total), accounting for 97.5% of the structure-aware learning task incorporated from DocStruct4M (Hu et al., 2024). TabMWP adds 38,431 tables (1.9% of the total); while smaller in scale, TabMWP offers valuable coverage of numerical reasoning examples.
@@ 3.3 Table Understanding Tasks
Tablet aggregates a total of 20 TU tasks grouped into 7 broad categories: Table Interpretation, Question Answering, Table-to-Text Generation, Table Numerical Reasoning, Table Natural Language Inference (NLI), and Table Structure Understanding. Subtasks within each category require different skills and demonstrate various degrees of complexity, ranging from basic table structure understanding to downstream tasks that combine tabular reasoning with information retrieval, numerical reasoning, or natural language inference. Table 1 provides a breakdown of examples per task and their source dataset. We briefly describe these below and provide more details in Appendix 7.5.
@@
Table Interpretation comprises three tasks, namely Column Type Annotation, Entity Linking, and Relation Extraction (ent_link, col_type, and rel_extraction in Table 1). Sourced from TURL (Deng et al., 2020), these tasks do not represent downstream applications but target basic table understanding skills essential for tackling more complex tasks. All instructions for these tasks were aggregated from TURL. They constitute the largest group in our training set, with 1,899,149 examples (54.2%).
@@
Table Question Answering is represented by Free-form Table QA (FeTaQA, Nan et al. 2022), Hierarchical Table QA (HiTabQA; Cheng et al. 2022), Hybrid QA (HybridQA; Chen et al. 2020b), and Table QA (WikiTableQuestions; Pasupat and Liang 2015). These tasks range from simple question answering over table content to multi-hop reasoning that combines textual and tabular information. They also vary in terms of table complexity and the expected answer format. Question answering represents 2.5% of Tablet’s training set (87,245 examples).
@@
Table-to-Text Generation is exemplified by two subtasks, namely Highlighted Cell Text Generation (ToTTo; Parikh et al. 2020) and Wikipedia Biography Generation (WikiBio; Lebret et al. 2016). Both tasks involve generating text based on table content, concise biographies in the case of WikiBio, and short descriptions based on visually highlighted cells, in the case of ToTTo. Together, they account for 693,593 training examples (19.8%).
@@
Table Numerical Reasoning combines Tabular Math Word Problem Solving (TabMWP; Lu et al. 2023) and Hybrid-context Financial QA (TAT-QA; Zhu et al. 2021). Given a table and a mathematical question, the model must perform reasoning over table values. This type of numerical reasoning represents 0.8% of Tablet’s training set (25,260 examples).
@@
Table NLI includes two entailment tasks: Table Fact Verification (TabFact; Chen et al. 2020a) and Infobox Natural Language Inference (InfoTabs; Gupta et al. 2020) where statements as supported or refuted based on table content. They represent 104,255 examples (3%) in Tablet’s training set.
@@
Table Structure Understanding is an umbrella category for tasks designed to facilitate structural understanding of tables: Merged Cell Detection, Row and Column Extraction, Table Cell Extraction, Table Cell Location, Table Recognition, Table Size Detection, and Structure-aware Parsing in Table 1). Examples for the first seven tasks are aggregated from MMTab, while Structure-aware Parsing is derived from DocStruct4M (Hu et al., 2024). These tasks are based on tables from multiple seed datasets (see rows with seeds M2, M1, and M3), including InfoTabs (Gupta et al., 2020), NSF (National Center for Science and Engineering Statistics, 2019), StatCan (Statistics Canada, 2024), TabMWP (Lu et al., 2023), TAT-QA (Zhu et al., 2021), TabFact (Chen et al., 2020a), ToTTo (Parikh et al., 2020), WikiBio (Lebret et al., 2016), WikiTableQuestions (Pasupat and Liang, 2015), and PubTabNet Zhong et al. (2020). While we maintain the original MMTab instructions, we use our own table visualizations. This group contributes 558,865 examples (15.8%) to our training.
@@
Visual Table Question Answering
Although many tasks in Tablet involve visually complex tables, they do not explicitly require models to exploit visual features beyond textual content. To evaluate whether training on Tablet’s original images improves models’ ability to integrate visual perception with table understanding, we introduce a manually curated Visual Table Question Answering benchmark. Human annotators selected tables with high visual complexity, as measured by our visual complexity scorer (Section 5 and Appendix 10), and formulated questions answerable only through joint reasoning over visual cues and tabular structure, e.g., ”What model is the red car in the table?” or ”In what year did the king holding a red cross die?”. To ensure questions genuinely require visual reasoning, we filtered out questions solvable from lossy synthetic representations alone. The resulting benchmark comprises 306 examples, each annotated with a question type to enable fine-grained performance analysis (see Appendix 13 for taxonomy and examples).